OpenStudy (akemskinner):

y = 5x – 1 –15x – 3y = 3 How many solutions does this linear system have? one solution: (0, –1) one solution: (1, 4) no solution infinite number of solutions

OpenStudy (mww):

both are linear so try to solve them simultaneously. These are the possibilities: If you get a sensible answer, there is one solution. If you answer does not include x and still makes sense, there are infinite solutions. (this occurs when the two equations you solve simultaneously are exactly the same) If your answer does not include x and does not make sense, there are no solution.

OpenStudy (akemskinner):

thanks

OpenStudy (mrnood):

y = 5x – 1 –15x – 3y = 3 re-arrange teh second equation: 3y =-3-15x divide by 3 y = - 5x-1 so now you have the following equations: y = - 5x-1 y = 5x – 1 see if oyu can solve those simultaneously